OpenStudy (anonymous):

Rewrite ^3 sqrt 27x-81 -5 to make it easy using a translation. Describe the graph. (The -5 is not square rooted)

OpenStudy (anonymous):

\[^{3}\sqrt{27x-81}-5\]

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Hmm, factor out the 27x-81 inside the radical

OpenStudy (dtan5457):

Getting a clue?

OpenStudy (dtan5457):

27x-81 27(x-3)

OpenStudy (anonymous):

Okay completely and entirely lost on this one.

OpenStudy (dtan5457):

\[\sqrt[3]{27x-81}-5=\sqrt[3]{27(x-3)}-5=3\sqrt[3]{x-3}-5\]

OpenStudy (dtan5457):

cube root of 27=3

OpenStudy (anonymous):

So is that it?

OpenStudy (dtan5457):

Yea

OpenStudy (anonymous):

Thank you!

OpenStudy (dtan5457):

No problem